ntegrate completely with respect to x :
∫21(4x+6)dx
1
Expert's answer
2014-09-11T13:08:42-0400
∫21(4x+6)dx=21∫(4x+6)dx=21∫4xdx+21∫6dx=84∫xdx+126∫dx=84x2/2+126x+C=42x2+126x+C, where C is an arbitrary real constant.